# General Chemistry Exam with Artificial Intelligence Jan Cvengros,a and Gerd Kortemeyerb†

### v1

@@ 1 Introduction
In many chemistry courses, assessment technologies remain anchored in closed-answer formats: multiple choice, short-answer numerical, or very short narrative responses 1. To probe argumentation and reasoning, instructors often must recast tasks into these closed formats to accommodate the technology 2. While there have been sustained efforts to embed higher-order thinking into introductory chemistry exams 3, assessment techniques such as oral examinations (one robust avenue for open-ended reasoning) and open-ended written responses are rarely scalable in large-enrollment settings 4. Yet open-ended problem solving is central to STEM assessment, offering rich visibility into conceptual understanding 5 and reasoning processes 6, 7. Compounding these tensions, the pandemic era amplified concerns about academic integrity in remote settings 8.
@@
Outside exam settings, frequent formative assessment further enhances learning through timely feedback 9, 10, but it is often underused because of grading workload. What students focus on is oftentimes driven by what is assessed and how much credit is assigned to the assessment, where ungraded exercises may be neglected 11. Traditional automated systems, generally limited to closed-format items (e.g., multiple choice or numeric responses), can yield strong psychometric reliability 12 but struggle to capture higher-order skills and can invite guessing 13. Longer answers, including creative exercises 14, mathematics-intensive exercises 15, drawing tasks 16, 17, and minute-papers 18, can foster formative self-reflection and help students in discovering knowledge structures and possible misconceptions.
@@
Against this backdrop, the rapid evolution of artificial intelligence, particularly large language models (LLMs) such as generative Pre-trained Transformers (GPT), is beginning to expand assessment options. Beyond scaling existing practices 19, recent work indicates that LLMs can evaluate complex, open-ended student responses at scale 20, 21, 22, 23, 24, 25. By reducing grading burden, these systems may make more frequent formative and low-stakes summative assessments feasible without overwhelming instructional resources 26, 27. In short, while current technologies have historically nudged assessments toward closed formats, advances in AI offer credible pathways to recover and systematically assess the richer reasoning and explanatory work that open-ended chemistry problems demand.
@@
Despite these potential benefits, integrating AI into grading workflows introduces critical considerations regarding grading reliability and transparency. AI-generated scores can occasionally exhibit undue confidence in ambiguous cases or misinterpret nuanced responses 28, 29, 30. Regulatory frameworks, such as the European Union’s designation of AI in educational assessments as high-risk, mandate human oversight to address these reliability concerns 31, 32, 33. Consequently, a human-in-the-loop grading approach is essential, ensuring that complex or ambiguous student responses receive necessary human attention 34.
@@
Grading provides valuable pedagogical benefits for teaching assistants (TAs), which may be an argument against using these systems. After all, by evaluating student work — particularly ambiguous or challenging cases — TAs deepen their own content knowledge and pedagogical understanding 35. AI-assisted grading systems can beneficially direct TA attention precisely toward these more instructive cases, filtering out routine grading tasks that offer limited educational value.
@@
This study investigates AI-assisted grading within a high-stakes, handwritten chemistry exam context. We employ psychometric methods, including Item Response Theory (IRT) 36, to evaluate the reliability of AI-generated scores, developing confidence thresholds that identify when human intervention is required. While our immediate focus is on high-stakes assessments, the methodologies presented are equally applicable to lower-stakes and purely formative scenarios, supporting the practical expansion of frequent assessments. Ultimately, this approach aims to balance efficiency with educational quality, promoting scalable, fair, and pedagogically rich assessment practices.
@@ 2.2 Exam
Fig.  1:  Examples of student work,. These would be uploaded as images to the model.
Fig.  2:  The rubric page associated with the student work in Fig. 1. This rubric image would be uploaded alongside the student work.
@@
  • The second problem focused on one hand on physical properties of liquids and gases, and students were asked to determine the boiling point, vapor pressure and enthalpy of vaporization of methanol based on graphical data. On the other hand, they analyzed systems, which reached thermal or chemical equilibrium. The former included the determination of the final temperature, the latter the calculation of reaction entropy, Gibbs free enthalpy of reaction or equilibrium constant as well as prediction of the effects of changing temperature and volume on chemical equilibria. The problem had 9 parts and was worth 13 points.
@@
  • Finally, the sixth problem tested the student’s understanding of the relationships between structure and properties. It also included a rather complex stoichiometric task related to mixing solutions of sulfuric acid with different concentrations. Furthermore, the students assessed different aspects of coordination compounds ($(\text{Ni(phen)}_{3})^{2+}$ and $(\text{Ni(H${}_{2}$O)}_{6})^{2+}$), including predicting molecular geometries, electron configurations or magnetic properties. They also examined their chirality and color based on ligand-field theory. The problem had 7 parts and was worth 1 points.
@@
Fig.  3:  Examples of the problem types on the exam.
All solutions were graded by teaching assistants (TAs) under the supervision of the instructor, who inspected all exams. A total of 16 teaching assistants were employed for five hours each.
@@ 3.1 Data collection
For students who chose to participate in our study, we scanned their 15-page exam packages into PDF using a standard copier/scanner with automated paper feed. We used a depersonalization scheme with numerical identifiers and a key that was only available to course personnel. This process took two individuals a little more than half a day. The pages of the PDFs were subsequently turned individual PNGs for each page and stored on a server. As it turned out, a handful of students had added a 16th sheet with additional explanations about certain problem parts; in accordance with the communicated exam policy, additional sheets were neither considered in the TA nor the AI grading of the exam..
@@ 3.2 AI systems used for grading
We used OpenAI models through Azure AI Services 37. The contractual framework stipulates that data is processed on European data centers and that data is not use for training purposes. In particular, we used the multimodal, reasoning model gpt-o4-mini/high-vision 38 (model version 2025-04-16).
@@
Fig. 4 shows the prompt used, which produced structured JSON output; this format is easy to process in subsequent steps. Each submission took between 50 and 60 seconds. At times we ran 30 agents in parallel, allowing us to grade the exams in about three hours. We used 32.15 million tokens, costing approximately \$100 in total; tokens are tiny chunks of text (roughly words or pieces of words) that the AI reads and writes, and because Azure charges per token, more text in prompts/context/responses means more tokens — and thus higher grading cost. This compares to $16\mbox{TAs}\times 5\mbox{h}\times\$44/\mbox{h}\approx\$3,500$ for grading the whole exam for all 459 students.
@@ 4. Return *only* a JSON array of objects, one per problem_part:
Fig.  4:  The prompt and JSON messages used for grading. The variables stu_url and rub_url would be replaced by the ephemeral URLs generated by our server.
@@
An important consideration in this workflow was to model a possible production setting as closely as possible. Grading assistance for an exam must not require in cumbersome data preparation or prohibitive amounts of manual steps. To broadly apply AI-grading assistance at scale, the net difference in effort must clearly be negative, and the process must be robust and intuitive enough to require minimal technical support.
@@ 3.3.1 Linear regression
A general quality measure for the AI-grading decisions is to perform linear regressions of the AI-assigned scores versus the established ground-truth scores. This approach assesses how closely the AI’s decisions align with human grading. The resulting linear regressions are characterized by three key parameters: slope, offset, and coefficient of determination (R?).
@@ 3.3.2 Classification performance metrics
With multiple classes $C_{i}$, the traditional measures for agreement of the model’s judgement and the ground truth are defined on a per-class base as
@@ 3.4 Confidence filters
Generative AI will always generate something. Besides the AI-grading itself, a second, equally important part is establishing confidence in the grading outcome 39. Confidence is different from quality measures, since in a production setting, the ground-truth would be unknown, and we need to establish confidence based on the AI-scores alone. For this, we turn to classical and Bayesian statistics.
Fig.  5:  Filtering AI-generated student-item grading results.

### v2

@@ 1.1 Open-ended responses
In many chemistry courses, assessment technologies remain anchored in closed-answer formats: multiple choice, short-answer numerical, or very short narrative responses mullen2012short. To probe argumentation and reasoning, instructors often must recast tasks into these closed formats to accommodate the technology stowe2019assessment. While there have been sustained efforts to embed higher-order thinking into introductory chemistry exams zoller2002algorithmic, assessment techniques such as oral examinations (one robust avenue for open-ended reasoning) and open-ended written responses are rarely scalable in large-enrollment settings gardner2023challenges. Yet open-ended problem solving is central to STEM assessment, offering rich visibility into conceptual understanding alsalmani23 and reasoning processes yik2021development, martin2023machine. Compounding these tensions, the pandemic era amplified concerns about academic integrity in remote settings clark2020testing.
@@
Outside exam settings, frequent formative assessment further enhances learning through timely feedback laverty12b, offerdahl2019formative, but it is often underused because of grading workload. What students focus on is oftentimes driven by what is assessed and how much credit is assigned to the assessment, where ungraded exercises may be neglected gomez2025student. Traditional automated systems, generally limited to closed-format items (e.g., multiple choice or numeric responses), can yield strong psychometric reliability pawl2013 but struggle to capture higher-order skills and can invite guessing towns2014guide. Longer answers, including creative exercises grieger2025utility, mathematics-intensive exercises ye2025just, drawing tasks baade2025seeing, stammes2025drawing, and minute-papers pentucci2025developing, can foster formative self-reflection and help students in discovering knowledge structures and possible misconceptions.
1.2 Automated grading
@@
Against this backdrop, the rapid evolution of artificial intelligence, particularly large language models (LLMs) such as generative Pre-trained Transformers (GPT), is beginning to expand assessment options. Beyond scaling existing practices lawrie2023establishing, recent work indicates that LLMs can evaluate complex, open-ended student responses at scale meyer2023chatgpt, kasneci2023chatgpt, wilson22, wan24, kortemeyer24aigrading, kortemeyer2024grading. By reducing grading burden, these systems may make more frequent formative and low-stakes summative assessments feasible without overwhelming instructional resources Floden2025, Fernandez2024. In short, while current technologies have historically nudged assessments toward closed formats, advances in AI offer credible pathways to recover and systematically assess the richer reasoning and explanatory work that open-ended chemistry problems demand.
@@
Despite these potential benefits, integrating AI into grading workflows introduces critical considerations regarding grading reliability and transparency. AI-generated scores can occasionally exhibit undue confidence in ambiguous cases or misinterpret nuanced responses ding2023students, dahlkemper23, li2023wrong. Regulatory frameworks, such as the European Union’s designation of AI in educational assessments as high-risk, mandate human oversight to address these reliability concerns euactannex, euactobl, de2020case. Consequently, a human-in-the-loop grading approach is essential, ensuring that complex or ambiguous student responses receive necessary human attention alsalmani23a, li2023learning.
@@
Grading provides valuable pedagogical benefits for teaching assistants (TAs), which may be an argument against using these systems. After all, by evaluating student work — particularly ambiguous or challenging cases — TAs deepen their own content knowledge and pedagogical understanding cho2010learning. AI-assisted grading systems can beneficially direct TA attention precisely toward these more instructive cases, filtering out routine grading tasks that offer limited educational value.
@@
This study investigates AI-assisted grading within a high-stakes, handwritten chemistry exam context. We employ psychometric methods, including Item Response Theory (IRT) rasch, to evaluate the reliability of AI-generated scores, developing confidence thresholds that identify when human intervention is required. While our immediate focus is on high-stakes assessments, the methodologies presented are equally applicable to lower-stakes and purely formative scenarios, supporting the practical expansion of frequent assessments. Ultimately, this approach aims to balance efficiency with educational quality, promoting scalable, fair, and pedagogically rich assessment practices.
1.3 Theoretical framework
We ground our approach in test theory and generalizability theory, treating AI-assigned points as fallible observations with multiple error sources (items, raters/runs, tasks) that can be estimated and, where appropriate, reduced through aggregation or design brennan1992generalizability. For the interpretation and use of AI-assisted scores, we adopt Kane’s argument-based validity perspective, which requires that claims about score meaning (and decisions such as pass/fail cut scores) be supported by explicit warrants and evidence kane2013validating. Our human-in-the-loop setup follows selective-prediction/deferral frameworks: the model grades confidently when reliable and defers uncertain cases to humans, analogous to the efforts in automated short-answer grading li2023learning.
1.4 Related efforts in AI-assisted grading
Within chemistry, early pilots have begun to use generative AI for automatic scoring of open-ended conceptual responses yamtinah2024_icce, sonkar2024automated, and complementary domain-specific approaches have automated evaluation of reaction answers within learning platforms plyer2024_softgrading. Across STEM, controlled studies report that large language models (LLMs) such as GPT-3.5/4 can score open-ended or short-answer exam responses with high agreement to human raters impey2025using, chen2025grading and improve further with fine-tuning and chain-of-thought prompting lee2024_patterns, latif2024_patterns. In higher-education settings, multi-course deployments have graded thousands of free-text answers (in multiple languages) using LLMs guided by instructor rubrics grevisse2024_bmcmeded, kortemeyer2025artificial, kortemeyer2024grading, kortemeyer2025assessing. In science education, human-in-the-loop workflows combine LLM scoring with explanations and selective review to increase trustworthiness cohn2024_cot. Benchmarking work aggregates diverse short-answer datasets and finds LLM-based graders outperform earlier baselines kortemeyer2023performance, yet still benefit from calibration and rubric alignment asag2024. Recent educational-data-mining studies add retrieval or multi-agent designs to stabilize grading on science short answers chu2025_edm_graderag, and multimodal pipelines now target scanned, diagram-rich or handwritten responses in science assessments zhu2025_ijcai. Together, these strands indicate a rapid shift toward AI-assisted scoring of open-ended work in chemistry and related STEM domains, while reinforcing the need for transparent reliability checks and human oversight.
@@ 2.2 Exam
Fig. 1: Examples of student work. These would be uploaded as images to the model. The exam was given in German, the translation of the problems is: Consider the pink-colored complex ion $\mbox{Ni}(\mbox{phen})_{3}]^{2+}$. The phen-ligand has the following structure: …Complete the following table: oxidation number of the metal, number of d-electrons of the metal ion, coordination number of the metal ion, number of ligands. Sketch the distribution of the d-electrons of the complex ion $\mbox{Ni}(\mbox{phen})_{3}]^{2+}$ in the crystal field of its ligands. Determine the magnetic properties of $\mbox{Ni}(\mbox{phen})_{3}]^{2+}$. Is the complex chiral? Provide the reasoning of your answer with an appropriate sketch. The student work on the right labelled the initial image of the structure, “2 free EP $\to$ bidentate ligand.” The student’s reasoning in the left panel is, “no, the two mirror images can be superimposed by rotation, even when the double bonds are taken into account;” the reasoning on the right is, after labelling the two sketches “mirror images,” “impossible to superimpose! $\Rightarrow$ it is chiral!”
Fig. 2: The rubric page associated with the student work in Fig. 1. The instructions are: +0.25 points for each identical answer to the table below. +1 point for following sketch (the arrows — number, distribution, and orientation — are most important; if the sketch is incorrect: +0.5 points for …(repeated for both configurations). +0.5 points for paramagnetic. +0.5 points for chiral; +0.5 points for the following sketch; +0.5 points for a mirror image of the above sketch.
@@
  • The second problem focused on the one hand on physical properties of liquids and gases, and students were asked to determine the boiling point, vapor pressure and enthalpy of vaporization of methanol based on graphical data. On the other hand, they analyzed systems, which reached thermal or chemical equilibrium. The former included the determination of the final temperature, the latter the calculation of reaction entropy, Gibbs free enthalpy of reaction or equilibrium constant as well as prediction of the effects of changing temperature and volume on chemical equilibria. The problem had 9 parts and was worth 13 points.
@@
  • Finally, the sixth problem tested the students’ understanding of the relationships between structure and properties. It also included a rather complex stoichiometric task related to mixing solutions of sulfuric acid with different concentrations. Furthermore, the students assessed different aspects of coordination compounds ($(\text{Ni(phen)}_{3})^{2+}$ and $(\text{Ni(H${}_{2}$O)}_{6})^{2+}$), including predicting molecular geometries, electron configurations or magnetic properties. They also examined their chirality and color based on ligand-field theory. The problem had 7 parts and was worth 10 points.
@@
Fig. 3: Examples of the problem types on the exam. For the drawing problem: draw the Lewis formula of the following particles with correct geometry (halfway correct binding angles and spatial structure including all free electron pairs and formal charges), and provide the structure type and molecular structure. For the graphing problem: draw the concentration of the anion $\mbox{SaI}^{2-}$ into the above diagram; the student wrote, “black line with slope +2, +1, and 0.” For the long answer: below you see the crystal structure of sodium hydride (light spheres: $\mbox{H}^{-}$, dark spheres $\mbox{Na}^{+}$). Describe the lattice arrangement with respect to the position of the individual ions;” the student wrote, “lattice structure reflects a typical NaCl type, the light spheres $\mbox{H}^{-}$ are located in the octahedral voids. On the other hand, the dark spheres, $\mbox{Na}^{+}$ cations, are cubic face-centered.” For the multiple-choice problem, the scenario is, “to an identical mixture like in problem part a), 100 ml of water are added. How do the parameters in the table below change? Mark the correct answer.” The choices are “decreases,” “stays the same,” and “increases” for each of the concentration of bromoethane, the overall reaction order, the rate constant, and the reaction rate. For the numerical problem: you have a sample of the nuclide $^{203}\mbox{Hg}$ with a mass of 20 mg. After how many days is only 1 mg of the original substance left? In the student response, “tagen” means days. For the reaction response: Permanganate ions $\mbox{MnO}_{4}^{-}$ oxidize hydrazine $\mbox{N}_{2}\mbox{H}_{4}$ to nitrogen in aqueous solution under basic conditions, forming manganese(IV) oxide. Formulate the stoichiometric reaction equation for this process. The short answer problem has already been discussed in Fig. 1. The symbolic problem reads, formulate the thermodynamic equilibrium constant $K$ for the formation of bromine chloride in accordance with the standard conventions.
@@ 2.3 Sample
2.4 Ground Truth
In this study, we treat the final human-assigned points as the ground truth, as it was produced by the course’s routine grading workflow. All solutions were graded by teaching assistants (TAs) under instructor supervision. To maximize within-item consistency, each problem or problem part was graded by a single responsible TA (“one-rater-per-item” design). Before independent scoring, the instructor and the TA co-graded approximately the first twenty solutions for that item, using a model solution with a detailed point allocation. An exceptions log documented adjudicated alternative solution paths and edge cases so that recurring situations received the same points; during the remainder of scoring, any novel cases were discussed with the instructor and added to the log. The instructor then conducted a brief audit of all graded exams. During the official post-exam review period, students could contest scores; disputed items were re-evaluated by the instructor and any corrections were entered before data export. However, only about one-third of the students took advantage of inspecting the grading, and less than a handful of incorrect grading decisions were discovered. In total, 16 TAs contributed five hours each (about 80 person-hours). Arguably, in spite of all due diligence, mistakes may have been made during grading and slipping through the control mechanisms, and arguably, one might refer to our ground truth as an operational reference standard, but it is that standard that AI-grading would held to.
@@ 3.1 Data collection
For students who chose to participate in our study, we scanned their 15-page exam packages into PDF using a standard copier/scanner with automated paper feed. While our contractual framework with the inference provider, Azure AI Services azure, is restrictive enough to permit processing personalized data, our IRB-approved research protocol did not allow this. We thus used a depersonalization scheme with numerical identifiers and a key that was only available to course personnel. The IRB protocol also stipulated that grading personnel were to be kept unaware which students signed the informed consent form to participate in the study, and which students did not.
This depersonalization and scanning process took two individuals a little more than half a day. The pages of the PDFs were subsequently turned into individual PNGs for each page and stored on a server. As it turned out, a handful of students had added a 16th sheet with additional explanations about certain problem parts; in accordance with the communicated exam policy, additional sheets were neither considered in the TA nor the AI grading of the exam.
@@ 3.2 AI systems used for grading
We used OpenAI models through Azure AI Services azure. The contractual framework stipulates that data is processed on European data centers and that data is not used for training purposes. In particular, we used the multimodal, reasoning model gpt-o4-mini/high-vision gpto4 (model version 2025-04-16).
@@
Fig. 4 shows the prompt used, which produced structured JSON output; this format is easy to process in subsequent steps. Each submission took between 50 and 60 seconds. At times we ran 30 agents in parallel, allowing us to grade the exams in about three hours. We used 32.15 million tokens, costing approximately \$100 in total; tokens are tiny chunks of text (roughly words or pieces of words) that the AI reads and writes, and because Azure charges per token, more text in prompts/context/responses means more tokens — and thus higher grading cost. This compares to $16\mbox{TAs}\times 5\mbox{h}\times\mathdollar 44/\mbox{h}\approx\mathdollar 3,500$ for grading the whole exam for all 459 students (USD 44.00 is the ETH-stipulated gross hourly income of CHF 30.70 plus employer contributions for pensions and insurance of about 15% at a 2025-exchange rate of $1.25$ USD/CHF).
@@ 4. Return *only* a JSON array of objects, one per problem_part:
Fig. 4: The prompt and JSON messages used for grading. The variables stu_url and rub_url would be replaced by the ephemeral URLs generated by our server.
@@
An important consideration in this workflow was to model a possible production setting as closely as possible. Grading assistance for an exam must not require cumbersome data preparation or prohibitive amounts of manual steps. To broadly apply AI-grading assistance at scale, the net difference in effort must clearly be negative, and the process must be robust and intuitive enough to require minimal technical support.
3.3 Inter-run reliability of points
[…]
Our primary metric is the intraclass correlation coefficient (ICC) shrout1979intraclass in its absolute-agreement, two-way random-effects form: ICC(A,1) for a single run and ICC(A,$k$) for the average of $k$ runs (here $k=5$). Conceptually, the ICC is the proportion of observed variation that reflects real differences between targets (students or cells) rather than run-to-run fluctuation. “Absolute agreement” means that any shift or scaling difference between runs counts against agreement; this is appropriate when exact points, not just rank order, matter. We estimate the ICC with the standard variance-components model (two-way random effects for targets and runs).
From the same model, we also report the within-target repeatability standard deviation $S_{w}$ (the typical run-to-run scatter for the same target) and the 95% repeatability coefficient $\mathrm{RC}=1.96\sqrt{2}\,S_{w}$. $\mathrm{RC}$ is directly interpretable on the point scale: for two independent runs, 95% of absolute differences for the same target are expected to be at or below $\mathrm{RC}$. In addition, because grading sometimes depends on ordering (e.g., relative rank), we summarize the stability of the ranking of total scores across runs using Kendall’s $W$ (0 to 1, with 1 indicating identical rankings) kendall1939problem. Finally, we quantify overall level agreement on totals with Lin’s concordance correlation coefficient (CCC) lawrence1989concordance, which combines correlation with penalties for any mean or scale bias between runs. For completeness, we also report the median and IQR of the per-student SD of total scores across runs as an intuitive measure of “wobble.”
All metrics are computed on the exam’s raw point scale, which ultimately determines grades (normalizing items by their maximum points is possible and would express results on a 0–1 scale, but we prioritize the raw scale because it aligns with grading decisions).
ICC(A,1) vs. ICC(A,$k$):
ICC(A,1) reflects reliability of a single AI run; ICC(A,$k$) shows how reliability would improve if one averaged $k$ runs (larger $k$ typically increases ICC).
$S_{w}$ and $\mathrm{RC}$ (points):
$S_{w}$ is the typical run-to-run SD for the same target. $\mathrm{RC}$ is a 95% bound on absolute differences between two runs; smaller values indicate tighter repeatability. Because it is on the point scale, $\mathrm{RC}$ can also be expressed as a percentage of the exam maximum.
Kendall’s $W$ (rank stability):
Values near 1 mean the ordering of students by total score is essentially unchanged across runs. Because, in many German-speaking university contexts, pass/fail thresholds or grade boundaries are set after inspecting the score distribution (rather than against a fixed absolute scale), stable rankings are particularly important.
Lin’s CCC (level agreement):
Values near 1 mean runs not only track each other strongly but also show negligible mean/scale bias (agreement along the identity line). Unlike Kendall’s $W$, CCC penalizes both mean and scale differences; for absolute point-scale bounds, see the repeatability metrics $S_{w}$ and $\mathrm{RC}$.
Per-student SD across runs:
The median SD (with IQR) gives an intuitive sense of how much individual totals vary across runs.
[…]
@@ 3.4.1 Linear regression
A general quality measure for the AI-grading decisions is to perform linear regressions of the AI-assigned scores versus the established ground-truth scores. This approach assesses how closely the AI’s decisions align with human grading. The resulting linear regressions are characterized by three key parameters: slope, offset, and coefficient of determination ($R^{2}$).
@@ 3.4.2 Classification performance metrics
With multiple classes $C_{i}$, the traditional measures for agreement of the model’s judgement and the ground truth are defined on a per-class basis as
@@ 3.5 Confidence filters
Generative AI will always generate something. Besides the AI-grading itself, a second, equally important part is establishing confidence in the grading outcome kortemeyer2025assessing. Confidence is different from quality measures, since in a production setting, the ground-truth would be unknown, and we need to establish confidence based on the AI-scores alone. For this, we turn to classical and Bayesian statistics.
Fig. 5: Filtering AI-generated student-item grading results.